OpenStudy (icalibear):

What is the common ratio of the sequence 80, 20, 5, 5/4... ?

OpenStudy (anonymous):

simplify 80 to 20 20 to 5 5 to 5/4

OpenStudy (icalibear):

How do I do that?

OpenStudy (anonymous):

kinda like simplifying the fraction 80/20 20/5 look for multiplies that they both share and divide those out

OpenStudy (icalibear):

20 is a common multiple for both would that work?

OpenStudy (anonymous):

for 80 to 20 yes

OpenStudy (icalibear):

so the common ratio is /20

OpenStudy (icalibear):

/ being division sign

hartnn (hartnn):

Common ratio = current term/previous term = next term / current term = 80/20 = 20/5 = 1/4 ok ?